Isayruiz:

In ABCD BH | AD and BG| CD if ab= 16 bc = 12 and bh= 14, find the length of altitude BG to side CD

Hero:

There should be a picture to with this question.

Hero:

The symbols you posted should likely be perpendicular symbols

Isayruiz:

Ok

1 attachment
Hero:

Okay give me a minute to take a look at this.

Hero:

So there are a few things we should understand regarding this particular problem. 1. Opposite sides and angles of parallelograms are congruent. 2. Understanding that, we need to show that \(\triangle ABH \sim \triangle{CBG}\) 3. Once we have shown that, we should be in good position to find what is required.

Isayruiz:

Okay I understand

Hero:

Basically, we know the triangles are similar because segments BH and AD are perpendicular. Segments BG and CD are perpendicular. Which means angles H and G are right angles.

Hero:

Furthermore, angles A and C are congruent. Because of this, we know triangles AHB and CGB are similar.

Hero:

Therefore, we can use the following proportion to find the length of BG: \(\dfrac{AB}{BH} = \dfrac{CB}{BG}\)

Isayruiz:

21/2 ?

Hero:

Okay so BG = 10.5. But now we need to find the length of segment CD. How might we figure out the length of segment CD?

Isayruiz:

Cd=16

Hero:

Very good. So what is the length of the altitude of segment BG to side CD?

Isayruiz:

10.5

Hero:

Well, actually it was asking for a fraction. The length of the altitude of segment BG would the numerator and the length of side CD would be the denominator.

Isayruiz:

21/2

Hero:

Not correct

Hero:

Hint: \(\dfrac{BG}{CD}\)

Hero:

BTW, We know 21/2 = BG. That goes in the numerator: \(\dfrac{21/2}{CD}\) You told me the length of \(\overline{CD}\) earlier so I know you know it.

Isayruiz:

What's the answer 21/32

Hero:

Correct

Isayruiz:

Thank you

Hero:

Welcome